OpenStudy (christos):

please help me find this limit \[\left\{ \frac{n!}{{{n}^{n}}} \right\}\underset{n=1}{\overset{\infty }{\mathop{{}}}}\,\]

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (christos):

qwll I have an idea but I am pretty sure that I am the only person in his world that is thinking that way

OpenStudy (christos):

but yes I solved it and its inf

OpenStudy (skullpatrol):

$$\Huge n!=1*2*3*4*...*n$$ $$\Huge n^n=\underbrace{n*n*n*...*n}$$ $$\Huge\text { _________n factors}$$

OpenStudy (christos):

\[n!=(n+1)(n+2)(n+3)...\]

OpenStudy (christos):

thats how I did it at least

OpenStudy (anonymous):

is it fractional part function?

OpenStudy (christos):

it seems you were right @skullpatrol indeed I found this at folmram https://www.dropbox.com/s/zq78o61ftgaweyf/Screenshot%202013-12-15%2015.16.12.jpg although it was unclear

OpenStudy (christos):

but I think both ways are right :D

OpenStudy (christos):

See? I told you I did it in a paradoxical way I am sure that I don't even know the real approach of these things

OpenStudy (skullpatrol):

Try to use the definitions as much as possible :)

OpenStudy (christos):

can you please tell me whats your approach ?

OpenStudy (christos):

@skullpatrol

OpenStudy (kc_kennylau):

http://math.info/image/290/ratio_test.gif

OpenStudy (kc_kennylau):

@Christos and \(n!\) is not \((n+1)(n+2)(n+3)\cdots\)

OpenStudy (kc_kennylau):

and it's not infinity

OpenStudy (christos):

dude I am trying to find the limit, no test I am allowed to use here

OpenStudy (christos):

no its infinity

OpenStudy (kc_kennylau):

ok sorry

OpenStudy (kc_kennylau):

but it's not infinity

OpenStudy (christos):

Why man?

OpenStudy (kc_kennylau):

why is it infinity

OpenStudy (christos):

that was my result

OpenStudy (christos):

whats yours?

OpenStudy (kc_kennylau):

my result is 0

OpenStudy (christos):

hm

OpenStudy (kc_kennylau):

can you show me your steps so that we can follow

OpenStudy (christos):

well let me type it in latex.........

OpenStudy (kc_kennylau):

ok :)

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}(n!/n ^{n})\]

OpenStudy (kc_kennylau):

lolz @nitz u don't say

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty}((n)(n-1)(n-2)........3.2.1)/(n.n.n..........n)\]

OpenStudy (kainui):

Well instead of thinking about the limit, let's look at something smaller but similar: 5!=1*2*3*4*5 5^5=5*5*5*5*5 In each case you're multiplying 5 numbers together, but the factorial is going to be multiplying smaller numbers while the exponent version is going to make all of the numbers the highest number that the factorial ever reaches.

OpenStudy (kc_kennylau):

lolz why don't we wait till she finishes typing in latex

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left\{ \frac{1*2*3...(n-1)*n}{n*n*n*n*n....{{n}^{n}}} \right\}_{n=1}^{\infty }\]

OpenStudy (kc_kennylau):

and the limit sign and the bracket are redundant

OpenStudy (christos):

I am a fast typer, just not used in MAthType

OpenStudy (kainui):

What's this (n=1) and (infinity) thing?

OpenStudy (kc_kennylau):

@Kainui that denotes a sequence

OpenStudy (christos):

Please tell mes wrong so that I fix it.

OpenStudy (kc_kennylau):

@Kainui and is misused in this scenario

OpenStudy (kainui):

It makes me think this is really some kind of power series. The difference between an infinite sum of these terms and just the limit of one of these terms is the difference between 0 and infinity.

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left[ \frac{1*2*3...(n-1)*n}{n*n*n*n*n....{{n}^{n}}} \right]_{n=1}^{\infty }\]

OpenStudy (kc_kennylau):

\[\Large\lim_{n\rightarrow\infty}\frac{n!}{n^n}\\\Large=\lim_{n\rightarrow\infty}\frac{1\times2\times\cdots\times n}{n\times n\times\cdots\times n}\\\Large=\lim_{n\rightarrow\infty}\frac1n\times\lim_{n\rightarrow\infty}\frac2n\times\cdots\times\lim_{n\rightarrow\infty}\frac nn\\\Large=0\times0\times\cdots\]

OpenStudy (kc_kennylau):

\[\Large=0\]

OpenStudy (christos):

http://screencast.com/t/W5N6M0kHu

OpenStudy (christos):

I think that's where you are mistaking

OpenStudy (kainui):

I think that's where you're mistaking @Christos.

OpenStudy (kc_kennylau):

the numerator is n! and the denominator is n^n

OpenStudy (christos):

yes I am talking about the denominator too.

OpenStudy (kainui):

Look, that entire part is n^n. Which is n times itself n times. Just like 3^3 is 3*3*3. and 4^4=4*4*4*4. It's n times itself n times.

OpenStudy (christos):

n*n*n*n*n..........n^n

OpenStudy (kainui):

No, wrong. n*n*n*n*...*n = n^n

OpenStudy (christos):

n*n*n*n*...*n but that kind of presentation implies that you know whats going on inside the "..."

OpenStudy (kc_kennylau):

http://screencast.com/t/NZAmQzMMvN

OpenStudy (christos):

forget about Kainui, I realised were I was wrong. I get it now.

OpenStudy (christos):

I see

OpenStudy (kc_kennylau):

you meant "forget about it Kainui"

OpenStudy (kainui):

No, the "..." represents that we don't know how many n's there are. 2^2=2*2 3^3=3*3*3 4^4=4*4*4*4 5^5=5*5*5*5*5 n^n=n*n*n*...*n Since n is a variable, we don't know how many times n is being multiplied by itself, so we can't possibly write out n n's!

OpenStudy (kainui):

Ok good =)

OpenStudy (christos):

yea yea I meant that I understood :D I missed the "it" haha

OpenStudy (kc_kennylau):

lolz those two have a great difference in meaning

OpenStudy (christos):

ironically yea :D look

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left[ \frac{{{n}^{n}}*n(cons\tan t)}{n} \right]_{n=1}^{\infty }\]

OpenStudy (kc_kennylau):

what's that constant

OpenStudy (christos):

wait

OpenStudy (kc_kennylau):

can u show me the original ques if it's on web

OpenStudy (christos):

the original questions? https://www.dropbox.com/s/6kpdq04ht4w19u6/Screenshot%202013-12-15%2015.52.14.jpg

OpenStudy (kc_kennylau):

which question is that?

OpenStudy (christos):

past 1. Question (ii)

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left[ \frac{{{n}^{n}}-n}{n} \right]_{n=1}^{\infty }\]

OpenStudy (kc_kennylau):

oh so there ain't a constant there lol (btw which question is that coz i can't find that in your list)

OpenStudy (christos):

That one: http://screencast.com/t/3VIez6jTgrop

OpenStudy (kc_kennylau):

I don't see any similarities between these: http://screencast.com/t/HIz71lvxv5W http://screencast.com/t/3VIez6jTgrop

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left[ n-1 \right]_{n=1}^{\infty }\]

OpenStudy (kc_kennylau):

come on don't type latex using the equation button it'll make you v......e.......r.....y...... slow

OpenStudy (christos):

I am uing MAthType program

OpenStudy (kc_kennylau):

if you can't use \(\LaTeX\), then \(\Huge\mbox{don't use it}\)

OpenStudy (christos):

I instead use latex, to can use it later

OpenStudy (kc_kennylau):

anyway learn to type \(\LaTeX\) as a code here

OpenStudy (christos):

:P

OpenStudy (christos):

hmm how?

OpenStudy (kc_kennylau):

\[\mbox{\[ \lim_{n\rightarrow\infty}\frac{n!}{n^n} \]}\]generates\[\lim_{n\rightarrow\infty}\frac{n!}{n^n}\]

OpenStudy (kc_kennylau):

don't bother with it if you can't use it, coz u can still survive with plain text: limit (n tends to infinity) n!/n^n

OpenStudy (christos):

I like latex :DDD

OpenStudy (kc_kennylau):

limit (n tends to infinity) n!/n^n =limit (n tends to infinity) (1*2*3*...*n)/(n*n*n*...*n) =limit (n tends to infinity) (1/n)*(2/n)*(3/n)*... (it won't end) =limit (n tends to infinity) 0*0*0*... =0

OpenStudy (christos):

So we get \[\left[ \infty -1 \right]-\left[ 1-1 \right]\]

OpenStudy (christos):

\[\left[ \infty -1 \right]-\left[ 1-1 \right]=\infty -0\]

OpenStudy (kc_kennylau):

can you show me how you derived from here: http://screencast.com/t/3VIez6jTgrop to here: http://screencast.com/t/HIz71lvxv5W

OpenStudy (christos):

ok sure

OpenStudy (skullpatrol):

you deserve a medal @kc_kennylau nice teaching

OpenStudy (christos):

\[\left\{ \frac{n!}{{{n}^{n}}} \right\}\underset{n=1}{\overset{\infty }{\mathop{{}}}}\,\] \[\underset{n\to \infty }{\mathop{\lim }}\,\left[ \frac{1*2*3*4...n(n-1)}{n*n*n...{{n}^{n}}} \right]_{n=1}^{\infty }\]

OpenStudy (christos):

I am sorry correction

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left[ \frac{1*2*3*4...n(n-1)}{n*n*n...n} \right]_{n=1}^{\infty }\] yes yes yes

OpenStudy (kc_kennylau):

stop putting the redundant brackets lol

OpenStudy (christos):

\[\underset{n\to \infty }{\mathop{\lim }}\,\left[ \frac{1*2*3*4...n}{n*n*n...n} \right]_{n=1}^{\infty }\] now it looks better :P

OpenStudy (experimentx):

use stirling's approximation